OpenStudy (anonymous):

Determine if the sequence is converging or diverging. If converging, find the limit. a(n)= (1+(-1)^n)/n Why is this converging?

OpenStudy (michele_laino):

please we have: \[{a_n} = \frac{2}{n}\] if n is even, and: \[{a_n} = 0\] if n is odd

OpenStudy (anonymous):

I'm sorry. I still don't know why.

OpenStudy (michele_laino):

when n is an even number then, we have: \[1 + {\left( { - 1} \right)^n} = 1 + 1 = 2\] whereas when n is an odd number, we have: \[1 + {\left( { - 1} \right)^n} = 1 - 1 = 0\]

OpenStudy (anonymous):

Yes, doesnt that indicate that it isn't monotonic?

OpenStudy (aum):

\[ a_n = \frac{1+(-1)^n}{n} = \frac 1n + \frac{(-1)^n}{n} \\ \text{As }n \rightarrow \infty, ~~~ \frac 1n \rightarrow 0 ~~~ \text {and }~~~ \frac{(-1)^n}{n} \rightarrow 0 \\ a_n \rightarrow 0 \text{ when } n \rightarrow \infty \\ \text{The series converges}. \]

OpenStudy (aum):

I meant the SEQUENCE converges. Not the series.

OpenStudy (anonymous):

Thanks. Sorry but I don't understand why (-1)^n / n = 0 as n -> infinity. I thought it was in some kind of oscillating path so I don't know why it is not DNE.

OpenStudy (aum):

(-1)^n / n If n is odd, numerator is -1. If n is even, numerator is +1. So the numerator is a finite constant (either -1 or +1). But the denominator is n which gets very large as n approaches infinity. So the quotient approaches 0.

OpenStudy (anonymous):

oh waittttt. oops. i get it now. thanks!

OpenStudy (aum):

If n = 1000, (-1)^n / n = 1 / 1000 = 0.001 If n = 1001, (-1)^n / n = -1 / 1001 approx = -0.001 Keep increasing n and you will find it approaches zero.

OpenStudy (aum):

You are welcome.